### v2

@@ 6.2. Ray tracing volumetric primitives
As described above, we limit the support of our kernels by bounding their bandwidth with an ellipsoid shell. We query the primitives along a ray leveraging ray-tracing, which we accelerate through the construction of a BVH using Optix (Parker et al., 2010) for efficient ray traversal. We implemented a ray-ellipsoid intersection as a custom intersection routine in Optix. However, we found that this produces suboptimal BVHs given the potentially large support of the primitives. Moreover, a custom intersection kernel is significantly less performant than a hardware-accelerated ray-triangle intersection test. Thus, instead of using the ray-ellipsoid intersection, we can triangulate the ellipsoids, and ray-trace on triangles. We analyzed a variety of triangulation approaches (see Figure 12), and found a significant boost on performance despite the higher number of primitives in the scene, finding highly-tesselated icospheres to perform best in terms of execution times. One could consider using mesh instancing to avoid duplicating the shell’s geometry. However, this approach reintroduces the problem of using axis-aligned bounding boxes for the individual instances in the instances BVH.

### v3

@@ 6.2. Ray tracing volumetric primitives
As described above, we limit the support of our kernels by bounding their bandwidth with an ellipsoid shell. We query the primitives along a ray leveraging ray-tracing, which we accelerate through the construction of a BVH using Optix (optix) for efficient ray traversal. We implemented a ray-ellipsoid intersection as a custom intersection routine in Optix. However, we found that this produces suboptimal BVHs given the potentially large support of the primitives. Moreover, a custom intersection kernel is significantly less performant than a hardware-accelerated ray-triangle intersection test. Thus, instead of using the ray-ellipsoid intersection, we can triangulate the ellipsoids, and ray-trace on triangles. We analyzed a variety of triangulation approaches (see Figure LABEL:fig:shell_ablation), and found a significant boost on performance despite the higher number of primitives in the scene, finding highly-tesselated icospheres to perform best in terms of execution times. One could consider using mesh instancing to avoid duplicating the shell’s geometry. However, this approach reintroduces the problem of using axis-aligned bounding boxes for the individual instances in the instances BVH.